OpenStudy (anonymous):

Solve the inequality: 3y-2<5/y

OpenStudy (anonymous):

the answer comes outs to 3y^2-2y-5=0 is this correct?

OpenStudy (anonymous):

\[ 3y^2-2y-5<0 \]Now you want to find to roots.

OpenStudy (anonymous):

Or rather, you'd find the roots to help factor. What you really want to do here is factor.

OpenStudy (anonymous):

answers come out to 5/3 and -1?

OpenStudy (anonymous):

Then \( (y-r_1)(y-r_2)<0 \) then \(y-r_1 < 0\) and \(y-r_2 > 0\) or the reverse.

OpenStudy (anonymous):

Since it has to be negative, one of them has to be negative, but not both and not none.

OpenStudy (anonymous):

so is it not 5/3 and -1

OpenStudy (anonymous):

Hold on, first of all we have: \[ \left(y-\frac 5 3\right)(y+1) < 0 \]Assuming your roots are true.

OpenStudy (anonymous):

We know that \[ \left(y-\frac 5 3\right) < (y+1) \]So that means if either one is going to be negative, then it will be \(y-5/3\)

OpenStudy (anonymous):

This gives us: \[ \left(y-\frac 5 3\right)<0 \]And \[ (y+1)>0 \]

OpenStudy (anonymous):

Or, simply put: \[ y < \frac 5 3, \,y>-1 \]

OpenStudy (anonymous):

Both must be true for the equality to be true.

OpenStudy (anonymous):

ok thanks